Hi..I was able to get part of the question, just not sure how to get the other part.
The earth as a radius of 6380 km. The formula for the volume of a sphere is V=4/3 pie r sq. The forula for surface area of a sphere is A= 4 pie r sq. Find the voume and surface area of the earth and write the answer in scientific notation with tw numbers after the decimal. The answer I got was 5.12 * 10 to the 8th power. Not sure how to get the second answer... thank you.

Answer by lwsshak3(11628)

Hi..I was able to get part of the question, just not sure how to get the other part.
The earth as a radius of 6380 km. The formula for the volume of a sphere is V=4/3 pie r sq. The forula for surface area of a sphere is A= 4 pie r sq. Find the voume and surface area of the earth and write the answer in scientific notation with tw numbers after the decimal. The answer I got was 5.12 * 10 to the 8th power. Not sure how to get the second answer... thank you.
...
Area of sphere:4πr^2
=4π(6380^2)
=4π*(6.38*10^3)^2
=4π*6.38^2*10^6
=511.5*10^6
=5.12*10^8 km^2
***
volume of a sphere is V=(4/3)πr^3
=(4/3)π(6380)^3
=(4/3)π*(6.38*10^3)^3
=(4/3)π*6.38^3*10^9
=1087.8*10^9
=10.88*10^11 km^3
**
ans:
Surface area of Earth=5.12 *10^8 km^2
Volume of Earth = 10.88*10^11 km^3